# An Information Theory of Finite Abstractions and their Fundamental Scalability Limits

### v1

@@ Abstract
Finite abstractions are discrete approximations of dynamical systems, such that the set of abstraction trajectories contains, in a formal sense, all system trajectories. There is a consensus that abstractions suffer from the curse of dimensionality: for the same “accuracy” (how closely the abstraction represents the system), the abstraction size scales poorly with system dimensions. And, yet, after decades of research on abstractions, there are no formal results concerning their accuracy-size tradeoff. In this work, we derive a statistical, quantitative theory of abstractions’ accuracy-size tradeoff and uncover fundamental limits on their scalability, through rate-distortion theory – the branch of information theory studying lossy compression. Abstractions are viewed as encoder-decoder pairs, encoding trajectories of dynamical systems in a higher-dimensional ambient space. Rate represents abstraction size, while distortion describes abstraction accuracy, defined as the spatial average deviation between abstract trajectories and system ones. We obtain a fundamental lower bound on the minimum abstraction distortion, given the system dynamics and a threshold on abstraction size. The bound depends on the complexity of the dynamics, through generalized entropy. We demonstrate the bound’s tightness on certain dynamical systems. Finally, we showcase how the developed theory can be employed to construct optimal abstractions, in terms of the size-accuracy tradeoff, through an example on a chaotic system.
@@ 1 Introduction
Modern engineering systems are becoming more complex and must meet intricate specifications in safety-critical situations. For instance, a self-driving car must follow traffic rules, avoid collisions, and optimize speed and fuel consumption. Due to the complexity of these systems, traditional analytic methods for verification and control are intractable. For over two decades, to address verification and control of complex dynamics and objectives, abstraction-based methods have flourished [1, 2]. Given a dynamical system, these methods construct a finite system – the abstraction –, arising from partitioning the state (and control) space of the original system, such that all trajectories of the original system are contained, in a formal sense, in the set of abstraction trajectories. Employing this property, one may solve an intractable verification or control problem for the original system over the finite abstraction, with formal guarantees of correctness. Over the years, research on abstractions has spanned deterministic systems [3, 4, 5], stochastic systems [6, 7, 8], and, lately, data-driven scenarios [9, 10, 11, 12].
Despite their immense success, there is a consensus that abstractions suffer from the curse of dimensionality, limiting their practical relevance; for a given accuracy (how closely the abstraction describes the true dynamics), the abstraction size scales poorly with system dimensions. And even though abstractions have received considerable interest in the past decades, there are still no formal results concerning their curse of dimensionality and accuracy-size tradeoff.
@@
In this work, we derive a statistical, quantitative theory of abstractions’ accuracy-size tradeoff and uncover fundamental limits on their scalability. To that end, we establish connections with rate-distortion theory – the branch of information theory studying lossy compression [13, Chapter 10]. The key observation for the whole theory is that abstractions are information-theoretic encoder-decoder pairs, encoding trajectories of dynamical systems, in a higher-dimensional, ambient space. Rate represents abstraction size, while distortion is defined as the spatial average deviation between abstract trajectories and system ones, thus capturing the average accuracy of an abstraction. Then, building on recent developments in rate-distortion theory for generalized measurable sets [14, 15], we derive fundamental limits of abstractions’ accuracy-size tradeoff: for given system dynamics, we obtain a fundamental lower bound on the minimum abstraction distortion, for a given threshold on abstraction size. The fundamental lower bound depends on the complexity of the system’s dynamics, through generalized entropy. We demonstrate the tightness of the bound on certain dynamical systems. Finally, we showcase how the developed theory can be employed to construct optimal abstractions, in terms of the size-accuracy tradeoff, through an example on a chaotic system, and we provide a discussion towards a general procedure for constructing optimal abstractions.
@@ 2.1 Measure spaces, Hausdorff measure, generalized entropy
For our purposes, we make use of information theory over general measurable spaces, based on [14, 15]. Thus, we first recall some related notions. We denote the $m$-dimensional Hausdorff measure [^2] The Hausdorff measure is a generalization of the Lebesgue measure, and measures the size of a given set. E.g., $\mathcal{H}^{1}(\mathcal{C})=2\pi$, where $\mathcal{C}$} is the unit circle embedded in $\mathbb{R}^{n}$. by $\mathcal{H}^{m}$. Denote the restriction of $\mathcal{H}^{m}$ to the compact set $\mathcal{K}$} by $\mathcal{H}^{m}_{{\mathcal{K}}}$. Consider a measure $\mu$ over the measure space $(\mathcal{X},\Sigma_{\mathcal{X}},\nu)$, where $\Sigma_{\mathcal{X}}$ is the Borel $\sigma$-algebra of $\mathcal{X}\subseteq\mathbb{R}^{n}$. When $\mu$ is absolutely continuous w.r.t. $\nu$ (denoted by $\mu\ll\nu$), we denote the Radon–Nikodym derivative by $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$. When $\nu=\mathcal{H}^{m}_{\mathcal{X}}$ (assuming $\mathcal{X}$ is $m$-dimensional), then $\frac{\mathrm{d}\mu}{\mathrm{d}\mathcal{H}^{m}_{\mathcal{X}}}$ is the probability distribution associated to $\mu$. Absolute continuity $\mu\ll\mathcal{H}^{m}_{\mathcal{X}}$ suggests that $\mu$ is not concentrated in arbitrarily small balls in $\mathcal{X}$. We denote the volume of the unit ball in $\mathbb{R}^{n}$ as $v_{n}=\frac{\pi^{n/2}}{\Gamma(n/2+1)}$, where the Gamma function $\Gamma(a)=\int_{0}^{\infty}t^{a-1}e^{-t}\mathrm{d}t$.
@@
The example below shows how the above apply to computing the entropy a dynamical system’s trajectories.
@@ 2.2 Rate-distortion theory on measurable spaces
A typical setting in information theory is source coding, see Fig. 1. A source emits a message $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$, which is a random variable over the measure space $(\mathcal{X},\Sigma_{\mathcal{X}},\mathcal{H}^{m}_{\mathcal{X}})$, with associated probability measure $\mu_{x}$, where $X$ is assumed to be $m$-dimensional. The encoder $s:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{Y}$ is finite, outputs the coded message $s(x)=y\in\mathcal{Y}$. Finally, the decoder $g:\mathcal{Y}\to\hat{\mathcal{X}}$, upon receiving $y$, decodes it into $g(y)=\hat{x}$. Compression takes place by encoding the continuous message $x$ into a low-dimensional, finite coded message $z$. The encoder cardinality $|\mathcal{Y}|$ determines the compression, and the compression rate is defined by $\log|\mathcal{Y}|$. A distortion function $d:\mathcal{X}\times\hat{\mathcal{X}}\to\mathbb{R}_{+}$ measures the deviation of $\hat{x}$ from the original message $x$. A typical distortion function, when $\hat{\mathcal{X}}=\mathbb{R}^{n}$, is the squared error $d(x,\hat{x})=\|x-\hat{x}\|^{2}$. [^3] Here, we present a simplified setting of source coding, where the encoder-decoder is deterministic, and the source emits a single message. For the general theory, see [13, 15].
@@
where the expectation is taken w.r.t. the random variable $x$. In words, $D(R)$ is the minimum achievable average distortion, for a given compression rate threshold $R$. The function $D(R)$ has an inverse, $R(D)$, which is the minimum compression rate, for a given maximum expected distortion threshold $D$. The following result provides a fundamental lower bound on $R(D)$ and $D(R)$.
@@
$R(D)$ $\geq R_{*}(D)\coloneqq h(x)-\frac{m}{2}-\log\Big(\frac{c_{\mathcal{X}}D^{m/2}\Gamma(1+m/2)}{(\frac{m}{2})^{m/2}}\Big),$ (3)
@@ 2.3 Transition systems
A transition system $S=(\mathcal{X},\underset{S}{\rightarrow})$ is deterministic if, for any $x\in\mathcal{X}$, there exists at most one $x^{\prime}\in\mathcal{X}$, such that $(x,x^{\prime})\in\underset{S}{\rightarrow}$. Given a transition system $S=(\mathcal{X},\underset{S}{\rightarrow})$, its $l$-length behavior ${\mathcal{B}}_{l}^{S}$ is defined as ${\mathcal{B}}_{l}^{S}:=\Big\{\xi:\ \xi=\{x_{i}\}_{i=0}^{l-1},\ (x_{k},x_{k+1})\in\underset{S}{\rightarrow},\ k=0,1,\dots,l-1\Big\}$. That is, the $l$-length behavior is the set of $l$-long trajectories. Notice that ${\mathcal{B}}_{l}^{S}\subseteq\mathcal{X}^{l}$.
@@ 3.1 Finite abstractions of dynamical systems
Although the dynamical system $S$ is deterministic, the abstraction AA is generally non-deterministic. With a slight abuse of formality, we often treat trajectories $\{\omega_{i}\}_{i=1}^{l}$ of the abstraction (with $\omega_{i}\in\mathcal{Y}$) as subsets of $\mathcal{X}^{l}$, that is $\{\omega_{i}\}_{i=1}^{l}\equiv\omega_{0}\times\omega_{1}\times\dots\times\omega_{l}\subseteq\mathcal{X}^{l}$.
@@
In fact, ${\mathcal{B}}_{l}^{A}$ is nlnl-dimensional, and covers the $n$-dimensional set of system trajectories ${\mathcal{B}}_{l}^{S}$. This observation is instrumental in this work. Through behavioral inclusion, abstractions encode information about the infinite, continuous system behavior ${\mathcal{B}}_{S}^{l}$ into the finite abstraction behavior set ${\mathcal{B}}_{l}^{A}$. While this enables computational methods to verification problems for dynamical systems, it also generally entails information loss, as the following section explains.
@@ 3.2 Abstraction-based verification and information loss
In typical verification problems, we are given a set of initial conditions $\Xi_{0}\subseteq\mathcal{X}$ for the system $S$ and we have to check if the corresponding set of system trajectories $\Xi=\{\xi\in{\mathcal{B}}_{l}^{S}:\ \xi_{0}\in\Xi_{0}\}$ satisfies a given property. For example, in the case of safety, we have to check if $\Xi\cap\mathcal{U}^{l}=\emptyset$, where $\mathcal{U}\subseteq\mathcal{X}$ is an unsafe set. Computing the exact reachable set $\Xi$ is generally impossible. Abstractions AA address this problem by computing the corresponding set of abstract state trajectories $\Omega_{A}=\bigcup\limits_{\omega\in{\mathcal{B}}_{l}^{A},\ \omega_{0}\cap\Xi_{0}\neq\emptyset}\omega$, which is tractable, as the abstraction is finite. Notice that, by behavioral inclusion, we have $\Xi\subseteq\Omega_{A}$. Finally, for safety verification, if $\Omega_{A}\cap\mathcal{U}^{l}=\emptyset$, then one may safely deduce that the system is safe.
As abstractions group system states $x\in\mathcal{X}$ in sets $Y\subseteq\mathcal{X}$, information loss is inevitable. In general, the partition $\mathcal{Y}$ needs to have a relatively high resolution, to recover a meaningful verification answer. E.g., in the extreme case of $|\mathcal{Y}|=1$, for any set of initial conditions $\Xi_{0}\subseteq\mathcal{X}$, the abstraction returns $\Omega_{A}=\mathcal{X}^{l}$, i.e. the whole ambient space of $l$-length trajectories. As such, for small $|\mathcal{Y}|$, the abstraction AA does not accurately represent the system $S$. On the other hand, for large $|\mathcal{Y}|$, where the abstraction is more accurate, the computations on the abstraction become heavier – even intractable. Thus, there is a trade-off between abstraction accuracy and partition size $|\mathcal{Y}|$. In what follows, we provide a statistical, quantitative theory of the accuracy-size tradeoff, based on rate-distortion theory, and provide bounds on the accuracy-size tradeoff.
@@ 4 Information-theoretic framework for finite abstractions
Let us show how an abstraction AA can be viewed as an encoder-decoder pair of system trajectories $\xi\in{\mathcal{B}}_{l}^{S}$. For the following, we refer the reader to Figure 2. The system (source) samples an initial condition $\xi_{0}\sim p_{\xi_{0}}$ and generates the trajectory $\xi=(\xi_{0},\dots,\xi_{l-1})\in{\mathcal{B}}_{l}^{S}$ (the message). The encoder $s_{A}:{\mathcal{B}}_{l}^{S}\to\mathcal{Y}$ looks at the initial condition $\xi_{0}$ and returns the corresponding abstract initial condition: [^5] Abstractions only use (sets of) initial conditions, for verification, as explained in Section 3. Nonetheless, from an information-theoretic perspective, $\xi_{0}$ and $\xi$ are equivalent, as $\xi_{0}\mapsto\xi$ is one-to-one; that is, $\xi_{0}$ and $\xi$ carry the exact same information when the dynamics $f$ is known.
@@
The compression rate, determined by the encoder’s size, is $\log(|\mathcal{Y}|)$. Indeed, notice that the abstraction encodes the system’s trajectories ${\mathcal{B}}_{l}^{S}$ into exactly $|\mathcal{Y}|$ outcomes, that is $\{g_{A}(z):\ z\in\mathcal{Y}\}$.
@@
In words, $d(\xi,\Omega_{A})$ returns the worst possible distortion between system trajectories $\xi$ and abstract trajectories $\Omega_{A}$, averaged over the time horizon $l$. This is in-line with abstraction-based verification, where the worst-case outcome is considered.
Let us now explain what “expected (or average) distortion”, for a given abstraction AA, means in the context of verification. The expected distortion $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$ is taken w.r.t. the initial-condition distribution $p_{\xi_{0}}$. Thus, for $N\to\infty$ verification problems, where the initial condition $\xi_{0}\sim p_{\xi_{0}}$, $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$ is the average distortion. As $d$ measures the distance between system trajectories and abstract state trajectories, the expected distortion $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$ is thus the spatial, statistical average of the deviation between system trajectories and abstract state trajectories, over initial conditions in $\mathcal{X}$ with distribution $p_{\xi_{0}}$.
@@
The distribution $p_{\xi_{0}}$ weights how much each initial condition $\xi_{0}\in\mathcal{X}$ contributes to the average distortion $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$. Arguably, the most suitable choice for $p_{\xi_{0}}$ is the uniform distribution, as, when constructing an abstraction, the initial condition is unknown and all initial conditions are considered equally likely.
@@
That is, the minimum average deviation of abstract state trajectories and system trajectories, over all possible abstractions with a given upper-bound $e^{R}$ on partition size. Likewise, we also consider the inverse $R_{abs}(D)$, which is the ($\log$ of the) minimum partition size for a given upper threshold $D$ on the average deviation of abstract state trajectories and system trajectories.

### v3

@@ Abstract
Finite abstractions are discrete approximations of dynamical systems, such that the set of abstraction trajectories contains all system trajectories. There is a consensus that abstractions suffer from the curse of dimensionality: for the same “accuracy” (how closely the abstraction represents the system), the abstraction size scales poorly with system dimensions. And yet, after decades of research on abstractions, there are no formal results on their accuracy-size tradeoff. In this work, we derive a statistical, quantitative theory of abstractions’ accuracy-size tradeoff and uncover fundamental limits on their scalability, through rate-distortion theory—the information theory of lossy compression. Abstractions are viewed as encoder-decoder pairs, encoding trajectories of dynamical systems. Rate measures abstraction size, while distortion describes accuracy, defined as the spatial average deviation between abstract trajectories and system ones. We obtain a fundamental lower bound on the minimum achievable abstraction distortion, given the system dynamics and the abstraction size; and vice-versa a lower bound on the minimum size, for given distortion. The bound depends on the complexity of the dynamics, through trajectory entropy. We demonstrate its tightness on some dynamical systems. Finally, we showcase how this new theory enables constructing minimal abstractions, optimizing the size-accuracy tradeoff, through an example on a chaotic system.
@@ 1 Introduction
Modern engineering systems are becoming more complex and must meet intricate specifications in safety-critical situations. For instance, a self-driving car must follow traffic rules, avoid collisions, and optimize speed and fuel consumption. Due to the complexity of these systems, traditional analytic methods for verification and control are intractable. For over two decades, to address verification and control of complex dynamics and objectives, abstraction-based methods have flourished [1, 2]. Given a dynamical system, these methods construct a finite system—the abstraction—, arising from partitioning the state (and control) space of the original system, such that all trajectories of the system are contained, in a formal sense, in the set of abstraction trajectories. Employing this property, one may solve an intractable verification or control problem for the original system over the finite abstraction, with formal guarantees of correctness. Over the years, research on abstractions has spanned deterministic systems [3, 4, 5], stochastic systems [6, 7, 8], and data-driven scenarios [9, 10, 11, 12].
Despite their immense success, there is a consensus that abstractions suffer from the curse of dimensionality, limiting their practical relevance; for a given accuracy (how closely the abstraction describes the true dynamics), the abstraction size scales poorly with system dimensions. And, as a rule of thumb, for better accuracy, a larger abstraction size is needed. However, after considerable interest on abstractions in the past decades, there are still no formal results concerning their curse of dimensionality and accuracy-size tradeoff.
@@
In this work, we derive a statistical, quantitative theory of abstractions’ accuracy-size tradeoff and uncover fundamental limits on their scalability. To that end, we establish connections with rate-distortion theory – the branch of information theory studying lossy compression [13, Chapter 10]. The key observation for the whole theory is that abstractions are information-theoretic encoder-decoder pairs, encoding trajectories of dynamical systems, in a higher-dimensional, ambient space. Rate represents abstraction size, while distortion is defined as the spatial average deviation between abstract trajectories and system ones, thus capturing the average accuracy of an abstraction. Then, building on recent developments in rate-distortion theory for generalized measurable sets [14, 15], we derive fundamental limits of abstractions’ accuracy-size tradeoff: we obtain a fundamental lower bound on the minimum achievable abstraction distortion, for given system dynamics and abstraction size; conversely, we also obtain a lower bound on the minimum abstraction size, for given distortion. The fundamental lower bound depends on the complexity of the dynamics, through generalized entropy, which we show how to compute. We demonstrate the tightness of the bound on certain dynamical systems. Finally, we showcase how the developed theory can be employed to construct optimal abstractions, in terms of the size-accuracy tradeoff, through an example on a chaotic system, and we provide a discussion towards a general procedure for constructing optimal abstractions.
@@ 2.1 Measure spaces, Hausdorff measure, generalized entropy
For our purposes, we make use of information theory over general measurable spaces, based on [14, 15]. Thus, we first recall some related notions. We denote the $m$-dimensional Hausdorff measure [^2] The Hausdorff measure is a generalization of the Lebesgue measure, generalizing the notions of curve length or surface area. E.g., $\mathcal{H}^{1}(\mathcal{C})=2\pi$, where $\mathcal{C}$} is the unit circle embedded in $\mathbb{R}^{n}$. by $\mathcal{H}^{m}$. Denote the restriction of $\mathcal{H}^{m}$ to the compact set $\mathcal{K}$} by $\mathcal{H}^{m}_{{\mathcal{K}}}$. Consider a measure $\mu$ over the measure space $(\mathcal{X},\Sigma_{\mathcal{X}},\nu)$, where $\Sigma_{\mathcal{X}}$ is the Borel $\sigma$-algebra of $\mathcal{X}\subseteq\mathbb{R}^{n}$. When $\mu$ is absolutely continuous w.r.t. $\nu$ (denoted by $\mu\ll\nu$), we denote the Radon–Nikodym derivative by $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$. When $\nu=\mathcal{H}^{m}_{\mathcal{X}}$ (assuming $\mathcal{X}$ is $m$-dimensional), then $\frac{\mathrm{d}\mu}{\mathrm{d}\mathcal{H}^{m}_{\mathcal{X}}}$ is the probability distribution associated to $\mu$. Absolute continuity $\mu\ll\mathcal{H}^{m}_{\mathcal{X}}$ suggests that $\mu$ is not concentrated in arbitrarily small balls in $\mathcal{X}$. We denote the volume of the unit ball in $\mathbb{R}^{n}$ as $v_{n}\coloneqq\frac{\pi^{n/2}}{\Gamma(n/2+1)}$, where the Gamma function $\Gamma(a)=\int_{0}^{\infty}t^{a-1}e^{-t}\mathrm{d}t$.
@@
The example below shows how the above apply to computing the entropy of a dynamical system’s trajectories.
@@ 2.2 Rate-distortion theory on measurable spaces
A typical setting in information theory is source coding, see Fig. 1. A source emits a message $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$, which is a random variable over the measure space $(\mathcal{X},\Sigma_{\mathcal{X}},\mathcal{H}^{m}_{\mathcal{X}})$, with associated probability measure $\mu_{x}$, where $\mathcal{X}$ is assumed to be $m$-dimensional. The encoder $s:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{Y}$ is finite, outputs the coded message $s(x)=y\in\mathcal{Y}$. Finally, the decoder $g:\mathcal{Y}\to\hat{\mathcal{X}}$, upon receiving $y$, decodes it into $g(y)=\hat{x}$. Compression takes place by encoding the continuous message $x$ into a low-dimensional, finite coded message $y$. The encoder cardinality $|\mathcal{Y}|$ determines the compression, and the rate is defined by $\log|\mathcal{Y}|$. A distortion function $d:\mathcal{X}\times\hat{\mathcal{X}}\to\mathbb{R}_{+}$ measures the deviation of $\hat{x}$ from the original message $x$. A typical distortion function, when $\hat{\mathcal{X}}=\mathbb{R}^{n}$, is the squared error $d(x,\hat{x})=\|x-\hat{x}\|^{2}$. [^3] Here, we present a simplified setting of source coding, where the encoder-decoder is deterministic, and the source emits a single message. For the general theory, see [13, 15].
@@
where the expectation is taken w.r.t. the random variable $x$. In words, $D(R)$ is the minimum achievable average distortion, for a given rate threshold $R$. The function $D(R)$ has an inverse, $R(D)$, which is the minimum rate, for a given maximum expected distortion threshold $D$. The following result provides a fundamental lower bound on $D(R)$.
@@ 2.3 Transition systems
A transition system $S=(\mathcal{X},\underset{S}{\rightarrow})$ is deterministic if, for any $x\in\mathcal{X}$, there exists at most one $x^{\prime}\in\mathcal{X}$, such that $(x,x^{\prime})\in\underset{S}{\rightarrow}$. Given a transition system $S=(\mathcal{X},\underset{S}{\rightarrow})$, its $l$-length behavior ${\mathcal{B}}_{l}^{S}$ is defined as ${\mathcal{B}}_{l}^{S}:=\Big\{\xi:\ \xi=\{x_{i}\}_{i=0}^{l-1},\ (x_{i},x_{i+1})\in\underset{S}{\rightarrow},\ i=0,1,\dots,l-1\Big\}$. That is, the $l$-length behavior is the set of $l$-long trajectories. Notice that ${\mathcal{B}}_{l}^{S}\subseteq\mathcal{X}^{l}$.
@@ 3.1 Finite abstractions of dynamical systems
Although the dynamical system $S$ is deterministic, the abstraction $A$ is generally non-deterministic. With a slight abuse of formality, we often treat trajectories $\{\omega_{i}\}_{i=1}^{l}$ of the abstraction (with $\omega_{i}\in\mathcal{Y}$) as subsets of $\mathcal{X}^{l}$, that is $\{\omega_{i}\}_{i=1}^{l}\equiv\omega_{0}\times\omega_{1}\times\dots\times\omega_{l}\subseteq\mathcal{X}^{l}$.
@@
In fact, ${\mathcal{B}}_{l}^{A}$ is $nl$-dimensional, and covers the $n$-dimensional set of system trajectories ${\mathcal{B}}_{l}^{S}$. This observation is instrumental in this work. Through behavioral inclusion, abstractions encode information about the infinite, continuous system behavior ${\mathcal{B}}_{S}^{l}$ into the finite abstraction behavior set ${\mathcal{B}}_{l}^{A}$. While this enables computational methods to verification problems for dynamical systems, it also generally entails information loss, as the following section explains.
Remark 1 (Extension to (approximate) simulations).
The results presented in this work straightforwardly extend to abstractions based on ($\varepsilon$-approximate) simulation relations (see [1]); for that, the abstraction state-space has to be a cover, instead of a partition, of $\mathcal{X}$. The corresponding set-based abstraction is obtained by picking $\varepsilon$ balls centered at the states of the point-based $\varepsilon$-approximate abstractions.
@@ 3.2 Abstraction-based verification and information loss
In typical verification problems, we are given a set of initial conditions $\Xi_{0}\subseteq\mathcal{X}$ for the system $S$ and we have to check if the corresponding set of system trajectories $\Xi=\{\xi\in{\mathcal{B}}_{l}^{S}:\ \xi_{0}\in\Xi_{0}\}$ satisfies a given property. For example, in the case of safety, we have to check if $\Xi\cap\mathcal{U}^{l}=\emptyset$, where $\mathcal{U}\subseteq\mathcal{X}$ is an unsafe set. Computing the exact reachable set $\Xi$ is generally impossible. Abstractions $A$ address this problem by computing the corresponding set of abstract state trajectories $\Omega_{A}=\bigcup\limits_{\omega\in{\mathcal{B}}_{l}^{A},\ \omega_{0}\cap\Xi_{0}\neq\emptyset}\omega$, which is tractable, as the abstraction is finite. Notice that, by behavioral inclusion, we have $\Xi\subseteq\Omega_{A}$. Finally, for safety verification, if $\Omega_{A}\cap\mathcal{U}^{l}=\emptyset$, then one may safely deduce that the system is safe.
As abstractions group system states $x\in\mathcal{X}$ in sets $Y\subseteq\mathcal{X}$, information loss is inevitable. In general, the partition $\mathcal{Y}$ needs to have a relatively high resolution, to recover a meaningful verification answer. E.g., in the extreme case of $|\mathcal{Y}|=1$, for any set of initial conditions $\Xi_{0}\subseteq\mathcal{X}$, the abstraction returns $\Omega_{A}=\mathcal{X}^{l}$, i.e. the whole ambient space of $l$-length trajectories. As such, for small $|\mathcal{Y}|$, the abstraction $A$ does not accurately represent the system $S$. On the other hand, for large $|\mathcal{Y}|$, where the abstraction is more accurate, the computations on the abstraction become heavier – even intractable. Thus, there is a trade-off between abstraction accuracy and partition size $|\mathcal{Y}|$. In what follows, we provide a statistical, quantitative theory of the accuracy-size tradeoff, based on rate-distortion theory, and provide bounds on the accuracy-size tradeoff.
@@ 4 Information-theoretic framework for finite abstractions
Let us show how an abstraction $A$ can be viewed as an encoder-decoder pair of system trajectories $\xi\in{\mathcal{B}}_{l}^{S}$. For the following, we refer the reader to Figure 2. The system (source) samples an initial condition $\xi_{0}\sim p_{\xi_{0}}$ and generates the trajectory $\xi=(\xi_{0},\dots,\xi_{l-1})\in{\mathcal{B}}_{l}^{S}$ (the message). The encoder $s_{A}:{\mathcal{B}}_{l}^{S}\to\mathcal{Y}$ looks at the initial condition $\xi_{0}$ and returns the corresponding abstract initial condition: [^5] Abstractions only use (sets of) initial conditions, for verification, as explained in Section 3. Nonetheless, from an information-theoretic perspective, $\xi_{0}$ and $\xi$ are equivalent, as $\xi_{0}\mapsto\xi$ is one-to-one, for given dynamics $f$; that is, $\xi_{0}$ and $\xi$ carry the exact same information.
@@
The rate, determined by the encoder’s size, is $\log(|\mathcal{Y}|)$. Indeed, notice that the abstraction encodes the system’s trajectories ${\mathcal{B}}_{l}^{S}$ into exactly $|\mathcal{Y}|$ outcomes, that is $\{g_{A}(z):\ z\in\mathcal{Y}\}$.
@@
In words, $d(\xi,\Omega_{A})$ is the worst possible distortion between system trajectories $\xi$ and abstract trajectories $\Omega_{A}$, averaged over the time horizon $l$. This is in-line with abstraction-based verification, where the worst-case outcome is considered.
Let us now explain what “expected (or average) distortion”, for a given abstraction $A$, means in the context of verification. The expected distortion $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$ is taken w.r.t. the initial-condition distribution $p_{\xi_{0}}$. Thus, for $N\to\infty$ verification problems, where the initial condition $\xi_{0}\sim p_{\xi_{0}}$, $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$ is the average distortion. As $d$ measures the distance between system trajectories and abstract state trajectories, the expected distortion $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$ is thus the spatial, statistical average of the deviation between system trajectories and abstract state trajectories, over initial conditions in $\mathcal{X}$ with distribution $p_{\xi_{0}}$.
@@
The distribution $p_{\xi_{0}}$ weighs how much each initial condition $\xi_{0}\in\mathcal{X}$ contributes to the average distortion $\mathbb{E}_{\xi_{0}}[d(\xi,\Omega_{A})\mid A]$. Arguably, the most suitable choice for $p_{\xi_{0}}$ is the uniform distribution, as, when constructing an abstraction, the initial condition is unknown and all initial conditions are considered equally likely.
@@
That is, the minimum average deviation of abstract state trajectories and system trajectories, over all abstractions with a given upper-bound $e^{R}$ on partition size. Likewise, we also consider the inverse $R_{abs}(D)$, which is the ($\log$ of the) minimum partition size for a given upper threshold $D$ on the average deviation of abstract state trajectories and system ones.